# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

### v1

@@ IV-F Few-shot Learning
Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain [62].
Definition IV.6
@@
(Few-shot Learning [62]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$.
@@
Few-shot learning is often characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1. Prior work primarily focused on applying meta-learning techniques to solve few-shot learning problems. These techniques overlap significantly with meta-learning for domain generalization. Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML [141], Prototypical network [172], Conditional Neural Processes [173])). Meta-learning techniques broadly fall into three categories: 1) parameter-based 2) metric-based 3) probabilitstic. Parameter-based meta-learning algorithms were discussed in Section III.G. An illustrative example demonstrating the primary difference between meta-learning for domain generalization and few-shot learning is provided in Figure 2.
@@
Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared [64, 174, 175]. Two neural networks are trained to encode data, from the source and target domain, into a latent feature space. One neural network, $g_{\theta}$, encodes the support set (i.e., source domain data), and a second, $f_{\theta}$, encodes the query set (i.e., target domain data). Typically, multiple support set data points are passed through the networks simultaneously to enable the calculation of a range of complex distance metrics.
(a) Conditional Neural Process
(a) Conditional Neural Process
(b) Latent Neural Process
Figure 3: A comparison of two basic neural process types: (a) conditional neural processes and (b) latent neural process. Green boxes represent input and output values, blue boxes represent a neural network, yellow boxes represent aggregated latent features, and the pink circle represent an aggregation function. This figure is adapted from Figure 3 in [176].
@@
Matching networks perform a pairwise comparison between latent features using cosine similarity [64]. Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space [172]. Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks [172]. Relation networks decompose the problem into separate modules: an embedding module and a relation module [177]. The embedding module is a neural network designed for feature extraction and the relation module is a non-linear classifier trained on the latent feature representation. Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class [178]. This technique maximizes the distance between subspaces by using a Grassmannian geometry-based objective function. Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g, [179, 180, 181]).
@@
Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process [182])) and a neural network. Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets [176]. Probabilistic meta-learning techniques fall into two broad categories: Neural Processes [183], and Deep Kernel Learning [184].
@@
Neural processes pass the data through a neural network, parameterized by $\theta$, to learn a latent representation. Each support set data point is passed through an aggregation function, $a$, to calculate the latent feature, $z$. This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes [176]). Another neural network, parameterized by $\phi$, uses $z$ and one or more data points (e.g., one-shot vs. few-shot) from the target domain to learn the parameters of a Gaussian Process. There are two main neural process types: latent neural processes, and conditional neural processes, as shown in Figure 3.
Latent neural processes construct a stochastic variable using latent features, which typically takes the form of a Gaussian distribution parameterized by $\mu$ and $\sigma$ [183]. However, by modeling these correlations the underlying predictive model becomes intractable and must be approximated (e.g., Monte Carlo Sampling [185], Variational Inference [183]). Conditional neural processes do not construct a stochastic variable using latent features [173], instead they make the assumption the stochastic process can be fully described by these encoded values (i.e., factorized Gaussian distribution [176]). This assumption makes the underlying predictive model analytically tractable, but ignores any correlations that may exist between latent features. Several neural network architectural components have been incorporated into conditional neural processes with varying degrees of success (e.g., [186, 187, 188, 189, 190, 191, 192, 193]).
@@
Deep kernel learning techniques takes an alternative by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, effectively training the network to learn a custom kernel function for that Gaussian process [184]. Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks [194]. This technique learns over a large number of small, but related, examples that train the underlying machine learning model to approximate a prior distribution to transfer knowledge between learning tasks. Deep kernel transfer quantifies uncertainty by maintaining a distribution over parameters of the neural network, which can become computationally expensive for larger networks. Pólya-Gamma augmentation and softmax approximation were explored to overcome this limitation by maintaining a distribution over functions [195]. Deep mean function learning has also been applied as a meta-learning technique, where a neural network is used to learn a custom mean function for a Gaussian process [196]. These and similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., [156, 197, 198, 199, 200]). However, recent results demonstrate that the original deep kernel learning algorithm tends to produce unreliable uncertainty estimates in some practical application domains [201]. Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (e.g., [201][202])
@@
Few-shot learning techniques exhibit many-to-many portability and high adaptability. Most few-shot learning techniques were developed using computer vision benchmarks (e.g., [203, 172, 64]), which requires deep learning architectures. However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model [11]. Therefore, it stands to reason that model complexity is directly related to the learning task’s complexity. Meta-learning techniques have been proven to be successful solutions to both few-shot learning and domain generalization. Further, these techniques exhibit desirable properties for all three presented criteria. Therefore, meta-learning presents a promising option for develop workload estimation for unknown tasks solutions and merit future investigation.
V Discussion
@@
Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are of particular importance to each task. Understanding the distribution across workload components is critical to adapting interactions and task allocations to maximize the performance of a human-robot team. Quantifying uncertainty helps enumerate this distribution by contextualizing component-specific workload estimates, as components may rely on overlapping metrics (e.g., fine-motor and tactile).
@@
Domain adaptation and transfer learning are popular methods, but require high model complexity to achieve desired performance levels. These methods are also best suited to one-to-one portability problems with low adaptability. Specifically, these techniques are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains and where all data is gathered prior to training. These techniques do not adapt machine learning models in the moment; thus, are not viable for performing workload estimation for unknown tasks in real world application domains. Test-time adaptation overcomes this particular constraint, but still exhibits one-to-one portability and low adaptability. Existing test-time adaptation techniques applied to smaller neural networks may exhibit higher adaptability, but further investigation is required.
@@
Continual learning has been frequently proposed as a potential solution to non-IID machine learning problems within HRI, but are not applicable to workload estimation for unknown tasks due to their poor portability. Incorporating information for an indeterminant number of tasks, while avoiding performance degradation on all prior tasks are additional constraints that likely make the machine learning problem unnecessarily difficult. These techniques learn how to transform the model between task pairs, whereas other techniques (e.g., domain generalization, few-shot learning) learn to transform from a common prior. Consider an application domain with $N$ tasks. Continual learning must learn $N!$ task-pair independent updates, as the tasks’ order influences the learning process. Domain generalization and few-shot learning techniques only learn $N$ updates, as they can trace each task to a single common prior. These additional portability-related constraints make continual learning a subpar option.
@@
Domain generalization and few-shot learning exhibit desirable properties for all three criteria: high adaptability, one-to-one portability, and low model complexity. These techniques are explicitly designed to update a machine learning model based upon limited and highly variable data, and are flexible enough change the model complexity based upon the learning task’s complexity. Multiple domain shifts from a common prior is the most directly related metaphor for the variability seen in human behavior; thus, domain generalization is the most suited for scenarios where there is no label shift (e.g., workload estimation). Few-shot learning may be more applicable to scenarios where labels change between learning tasks or application domains (i.e., task recognition, affect recognition). The presented meta-learning techniques are general enough to apply to either machine learning problem.
@@
Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem [1, 204, 21]. Only a limited number of techniques (e.g., probablistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks. Probablistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
@@
A key research question is if existing workload estimation datasets capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques can accurately estimate workload for unknown tasks. It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial. These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID and non-IID tasks that these techniques can handle. Understanding how probablistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real world viability. Additionally, analyzing how the differences between tasks, human-robot interactions, robot capabilities further informs how these techniques can be successfully deployed. Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown tasks; thus, merit further investigation.
VI Conclusion
@@
Developing models that can estimate the human’s internal state is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments. The variability of these environments is characterized by varying human-robot teaming dynamics, diverse tasks executions, and different operating conditions for the team itself. A robot seeking to adapt its behavior to best assist its human teammate must accurate the human’s internal state accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible. Non-IID machine learning, particular domain generalization and few-shot learning, hold promise in developing models that can account for these varying dynamics.
Acknowledgments
The graduate students have been supported by National Aeronautics and Space Administration through University Space Research Association aware 904186092. The contents are those of the authors and do not represent the official views of, nor an endorsement, by the National Aeronautics and Space Administration.
The presented work was partially supported by ONR grants N00024-20-F-8705 and N00014-21-1-2052. The views, opinions, and findings expressed are those of the authors and are not to be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.

### v2

@@ 4.6 Few-shot Learning
Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain (Wang et al., 2020b).
Definition:.
@@
(Few-shot Learning) Given a set of source domains $DS$ = $D_{s_{1}},\dots,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},\dots,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$ (Wang et al., 2020b).
Few-shot learning is characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1. This problem formulation is similar to domain generalization, but with additional complexity. A domain shift results from the novel physiological response that an unknown task can induce, but a label shift is a new workload value. An illustrative example demonstrating the primary difference is provided in Figure 2. The left task has high auditory and fine motor workload, but a low gross motor workload. The right task has a high gross motor workload, but a low auditory workload. Training a model to estimate gross motor workload on the left task will result in a model that likely performs poorly when estimating workload for the right task, as the range of gross motor workload values is too different. Few-shot learning techniques overcome the additional complexity of handling multiple distribution shifts simultaneously by requiring labeled target domain data.
@@
Prior work primarily focused on applying meta-learning techniques to solve few-shot learning problems. These techniques overlap significantly with meta-learning for domain generalization. Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML (Finn et al., 2017), Prototypical network (Snell et al., 2017), Conditional Neural Processes (Garnelo et al., 2018a))). Meta-learning techniques broadly fall into three categories: 1) parameter-based, 2) metric-based, and 3) probabilistic. Parameter-based meta-learning algorithms were discussed in the Domain Generalization section.
@@
Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared (Vinyals et al., 2016; Zhou et al., 2023a; Li et al., 2023). Two neural networks are trained to encode data, from the source and target domain, into a latent feature space. One neural network, $g_{\theta}$, encodes the support set (i.e., source domain data), and a second, $f_{\theta}$, encodes the query set (i.e., target domain data). Typically, multiple support set data points are passed through the networks simultaneously to enable the calculation of a range of complex distance metrics.
@@
Matching networks perform a pairwise comparison between latent features using cosine similarity (Vinyals et al., 2016). Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space (Snell et al., 2017). Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks (Snell et al., 2017). Relation networks decompose the problem into separate modules: an embedding module and a relation module (Sung et al., 2018). The embedding module is a neural network designed for feature extraction and the relation module is a non-linear classifier trained on the latent feature representation. Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class (Simon et al., 2020). This technique maximizes the distance between subspaces by using a Grassmannian geometry-based objective function. Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g., Kim et al., 2019; Schonfeld et al., 2019; Zhang et al., 2019).
@@
Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process (Rasmussen, 2003)) and a neural network. Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets (Jha et al., 2022). Probabilistic meta-learning techniques fall into two broad categories: Neural Processes (Garnelo et al., 2018b), and Deep Kernel Learning (Wilson et al., 2016).
@@
Neural processes pass the data through a neural network to learn a latent representation. Each support set data point is passed through an aggregation function to calculate the latent feature. This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes (Jha et al., 2022)). Another neural network uses that latent representation and one or more data points (e.g., one-shot vs. few-shot) from the target domain to learn the parameters of a Gaussian Process. There are two main neural process types: latent neural processes (Garnelo et al., 2018b) and conditional neural processes (Garnelo et al., 2018a). Conditional neural processes have received the most attention, where many techniques incorporate different neural network architectural components with varying degrees of success (e.g., Gu et al., 2023a; Holderrieth et al., 2021; Bruinsma et al., 2023; Kim et al., 2022; Wang et al., 2022; Ye and Yao, 2022; Dexheimer and Davison, 2023; Pandey and Yu, 2023)).
@@
Deep kernel learning techniques take an alternative approach by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, training the network to learn a custom kernel function for that Gaussian process (Wilson et al., 2016). Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks (Patacchiola et al., 2020). This technique learns over a large number of small, but related, examples that train the machine learning model to approximate a prior distribution to transfer knowledge between learning tasks. Deep kernel transfer quantifies uncertainty by maintaining a distribution over parameters of the neural network, which can become computationally expensive for larger networks. Similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., Rothfuss et al., 2023; Wei et al., 2023; Bånkestad et al., 2023; Maraval et al., 2024; Iwata and Kumagai, 2024; Lee et al., 2023). However, recent results demonstrate that many deep kernel learning algorithms tend to produce unreliable uncertainty estimates in some application domains (van Amersfoort et al., 2021). Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (Liu et al., 2023).
@@
Few-shot learning techniques exhibit many-to-many portability and high adaptability. Most few-shot learning techniques were developed using computer vision benchmarks (e.g., Garcia and Bruna, 2018; Snell et al., 2017; Vinyals et al., 2016), which requires deep learning architectures. However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model (Tian et al., 2020). Therefore, it stands to reason that model complexity is directly related to the learning task’s complexity. Meta-learning techniques have been proven to be successful solutions to both few-shot learning and domain generalization. Further, these techniques exhibit desirable properties for all three presented criteria. Therefore, meta-learning presents a promising option for develop workload estimation for unknown tasks solutions and merit future investigation.
5 Discussion
@@
Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are particularly important for each task. Understanding the distribution across workload components is critical to adapting robot behavior to maximize the human-robot team’s performance.
Quantifying uncertainty helps enumerate this distribution by contextualizing component-specific workload estimates, as components may rely on overlapping metrics (e.g., fine-motor and tactile). Viewing candidate machine learning methods through the three developed criteria (i.e., portability, model complexity, and adaptability) facilitates discussing for how each candidate addresses the core concerns of estimating human workload for unknown tasks (see Evaluation Criteria section).
@@
Domain adaptation and transfer learning are popular methods, but are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains, and where all data is gathered prior to training (i.e., one-to-one portability and low adaptability). These methods are not well suited for estimating a human’s workload for unknown tasks, as it is impractical to gather a significant amount of data for new tasks (i.e., target domain) prior to training the model. Additionally, these techniques do not adapt machine learning models in real time. Test-time adaptation overcomes this particular constraint, but exhibits one-to-one portability and low adaptability. Existing test-time adaptation techniques applied to smaller neural networks may exhibit higher adaptability, but further investigation is required. Regardless, none of these techniques are well suited for estimating human workload in the presence of unknown tasks.
@@
Continual learning has been frequently proposed as a potential solution to non-IID machine learning problems within HRI, but are not applicable to workload estimation for unknown tasks due to their poor portability. Incorporating information for an indeterminant number of tasks, while avoiding performance degradation on all prior tasks are additional constraints that likely make the machine learning problem unnecessarily difficult. These techniques learn how to transform the model between task pairs, whereas other techniques (e.g., domain generalization, few-shot learning) learn to transform from a common prior. Consider an application domain with $N$ tasks. Continual learning must learn $N!$ task-pair independent updates, as the tasks’ ordering influences the learning process. Domain generalization and few-shot learning techniques only learn $N$ updates, as they can trace each task to a single common prior. These additional portability-related constraints make continual learning a subpar option.
@@
Domain generalization and few-shot learning exhibit desirable properties for all criteria: high adaptability, one-to-one portability, and low model complexity. These techniques are designed to update a machine learning model based on limited and highly variable data, and are flexible enough to change the model complexity based on the learning task’s complexity. Multiple domain shifts from a common prior is the most directly related metaphor for task-based variability, as the contribution of each physiological signal to each workload component and the contribution of each workload component will vary between tasks. Therefore, domain generalization is the most suited for workload estimation for unknown tasks when there is no label shift. Few-shot learning is more applicable to scenarios where there are significant changes to the range of workload values between tasks (i.e., label shift). Meta-learning techniques sit at the intersection of these two categories and are general enough to apply to either learning task.
@@
Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem (Heard et al., 2019; Bhagat Smith et al., 2023, 2022). Only a limited number of techniques (e.g., probabilistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks. Probabilistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
@@
Existing workload estimation datasets frequently struggle to capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques’ ability to accurately estimate workload for novel tasks can be assessed. It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial (Bhagat Smith et al., 2024). These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID (i.e. known) and non-IID tasks (i.e., unknown) that these techniques can handle. Understanding how probabilistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real-world viability. Additionally, analyzing how the differences between tasks, human-robot interactions, and robot capabilities further informs how these techniques can be successfully deployed. Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown novel tasks; thus, merit further investigation.
6 Conclusion
@@
Developing models that can estimate the human’s workload is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments. The variability of these environments is characterized by varying human-robot teaming dynamics, diverse tasks executions, and different operating conditions for the team itself. A robot seeking to adapt its behavior to best assist its human teammate must estimate the human’s workload accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible. This manuscript developed three criteria (i.e., portability, model complexity, and adaptability) and presented an analysis of relevant non-IID machine learning techniques. Numerous non-IID machine learning techniques exist, so this manuscript assessed how each technique’s machine learning considerations interact with the unique characteristics of HRI datasets (e.g., workload estimation for unknown tasks). A deeper understanding of these interactions highlight the specific requirements for a particular application, which is useful for guiding future empirical research. This manuscript’s analysis argues that domain generalization and few-shot learning techniques hold promise in developing models that can estimate a human’s workload for unknown tasks.